# X-ray Emission Signatures of Neutron Star Mergers

### v1

@@ 4.1.2 BNS-II (BNS $\rightarrow$ HMNS $\rightarrow$ SMNS $\rightarrow$ BH)
  • Jet zone: Bright prompt emission and decaying afterglow from the GRB jet. The extended wind emission becomes observable once the prompt emission sufficiently diminishes, and will suddenly drop when the remnant collapses. See left panel of Fig. 9.
  • Free zone: Faint prompt emission and rising afterglow from the GRB jet. The extended X-ray wind emission becomes observable once the prompt emission sufficiently diminishes—in some cases the prompt emission will never exceed the wind emission and only a plateau will be observed. The plateau will end abruptly when the remnant collapses. See middle panel of Fig. 9.
  • Trapped zone: Extremely faint prompt emission and dim late-time rising afterglow from the GRB jet. If $t_{\rm thin}<t_{c}$, the extended wind emission will become directly observable. See right panel of Fig. 9.
@@
4.1.3 BNS-III (BNS $\rightarrow$ HMNS $\rightarrow$ BH)
Physically, one would expect only prompt X-rays from the GRB jet. Observationally, in X-rays, one would observe various emission signatures depending on the viewing angle,
@@ 4.1.3 BNS-III (BNS $\rightarrow$ HMNS $\rightarrow$ BH)
4.1.4 BNS-IV (BNS $\rightarrow$ BH)
Remnants with masses much larger than the TOV mass limit will be unable to withstand collapse into a BH for an appreciable amount of time. A jet will still be launched and the associated light curve will likely resemble a BNS-III (Fig. 10)
4.1.5 The case of GW 170817
The EM counterparts of GW 170817, including both prompt and afterglow emission, can be consistently interpreted within the framework presented here. However, since the earliest X-ray detections of GW 170817 were not until hours after the initial trigger, it remains unclear whether the NS merger of GW 170817 was a BNS-II or BNS-III/BNS-IV. Prompt X-ray detections of NS mergers by a wide-field X-ray detector such as EP-WXT can identify the presence of a long-lived central engine and distinguish between these scenarios.
4.2 BH-NS Mergers
X-ray counterparts are expected only from BH-NS mergers that tidally disrupt the NS (BH-NS w/ TD). Physically, one would expect prompt and afterglow X-rays from a GRB jet. Compared to jets launched from BNS mergers, jets launched from BH-NS mergers w/ TD are likely less collimated (Bromberg et al., 2011; Duffell et al., 2015; Ruiz et al., 2018; Lazzati & Perna, 2019). Additionally, the dynamical ejecta should be directed primarily in the equatorial direction (Kawaguchi et al., 2015, 2016). Therefore, we adopt $\theta_{c}=15^{\circ}$ and $\theta_{\rm cut}=60^{\circ}$ for a BH-NS w/ TD jet and adjust $E_{\rm iso}$ such that the beaming-corrected energy $E_{\gamma}=(1-\cos\theta_{c})\,E_{\rm iso}$ remains constant. The BH-NS w/ TD jet, while less energetic than its BNS counterpart, has a shallower energy and Lorentz factor profile and larger cutoff angle, thus appearing brighter at large viewing angles.
@@ 4.2 BH-NS Mergers
4.3 X-ray Counterpart Searches to Neutron Star Mergers
The search for EM counterparts of NS mergers detected by the LIGO-Virgo-KAGRA collaboration (LVK) is critical for advancing understanding of both the merger process and the Type I GRBs. The joint detection of GW170817 and GRB 170817A has opened a new era of multi-messenger astronomy, demonstrating the power of combining GW and EM observations.
@@ 4.3 X-ray Counterpart Searches to Neutron Star Mergers
The newest generation of wide-field X-ray telescopes, including the Einstein Probe (EP; Yuan et al. (2022)) and the Space-based multi-band astronomical Variable Objects Monitor (SVOM; Wei et al. (2016)), boasts rapid response and dedicated sensitivity in the soft X-rays. EP’s Wide-field X-ray Telescope (WXT), with its wide field-of-view and high sensitivity in the soft X-ray band (0.5–4 keV), is designed to promptly detect and localize soft X-ray transients across broad sky regions. Its on-board Follow-up X-ray Telescope (FXT), is a very sensitive X-ray instrument that allows deep follow-up observations in X-rays. SVOM offers complementary multi-wavelength capabilities through its diverse payload: the wide-field ECLAIR ($4-150\rm\,keV$) enables prompt detections, while the Microchannel X-ray Telescope (MXT; $0.2-10\rm\,keV$) facilitates rapid X-ray follow-up. Additionally, SVOM is equipped with rapid slewing capability and is integrated with ground-based follow-up networks, enhancing its effectiveness in capturing transient events.
@@
Fig. 6 shows the X-ray peak luminosity of prompt emission from NS mergers with EM counterparts observed at different viewing angles, assuming an isotropic-equivalent energy in gamma-rays $E_{\rm iso}=10^{51}$ erg. For BNS mergers, the jet opening angle is typically narrow, with $\theta_{c}\sim 5^{\circ}$, as observed in GRB 170817A (Troja et al., 2017, 2019b, 2020), whereas for BH-NS mergers, the opening angle is likely larger, $\theta_{c}\sim 15^{\circ}$ (Bromberg et al., 2011; Duffell et al., 2015; Ruiz et al., 2018; Lazzati & Perna, 2019). Therefore, early prompt X-ray detections—or non-detections—with luminosity distance information inferred from LVK events could place meaningful constraints on the viewing angles of these mergers.
Figure 12: X-ray afterglow peak luminosity (left panel) and peak time (middle panel) as functions of viewing angle for a GW170817-like event. The right panel shows the correlation between peak luminosity and peak time. The beaming-corrected energy is fixed at $10^{51}$ erg. X-ray data was retrieved from Troja et al. (2020).
@@
With this constraint, one may calculate the X-ray afterglow peak time and luminosity for reasonable Type I GRB parameters. Figure 12 shows the X-ray afterglow peak luminosity (Left panel) and peak time (Middle panel) as functions of the viewing angle $\theta_{v}$. The parameters are listed in Table 2. With the inferred luminosity distance from LVK, observers may estimate the detection horizon for X-ray counterparts of NS mergers based on their detector’s sensitivity and time since the GW trigger (Fig. 14).
Figure 14: Detection horizon for X-ray (0.3-10 keV) afterglows from BNS/BH-NS mergers as a function of time, for given detector sensitivities. The model parameters are listed in Tab. 2.
@@
  • Prompt X-ray Analysis: Data from X-ray detectors that were covering the localized region during the trigger time should be examined for coincident events. If no events are found, upper limits can help constrain the prompt X-ray emission.
  • Viewing Angle Constraints: Constraints can be obtained by X-ray detections and upper limits, using the model described in Section 3.1 and illustrated in Fig. 6. The inferred isotropic-equivalent luminosity, $L_{\text{iso}}=E_{\text{iso}}/T_{90}$, constrains the viewing angle $\theta_{v}$.
  • X-ray Afterglow Predictions: With constraints on $\theta_{v}$, the X-ray afterglow characteristics can be estimated, e.g. the peak time and flux (Fig. 12).
@@
  • Coordinated Follow-up: X-ray follow-up observations can be planned based on the predicted afterglow properties. Using Figure 14, one can estimate the optimal time window for detecting an X-ray counterpart of an NS merger with known luminosity distance, given one’s detector sensitivity. It suggests that prompt follow-up observations are very important. In case of non-detection at a certain epoch, one can estimate the luminosity distance horizon beyond which continued follow-up is no longer necessary.
Additional detections of NS mergers with X-ray counterparts will be a major success of multi-messenger astronomy, and rapid follow-up observations play a crucial role in enabling such discoveries. According to Figure 14, if EP-FXT (with sensitivity of about $10^{-14}\ {\rm erg/s/cm^{2}}$) does not detect an X-ray counterpart of a BH-NS merger within one day, then follow-up observations are no longer needed if the source is further than $\sim 50$ Mpc. For BNS mergers, EP-FXT follow-up observations after one day are no longer needed for sources beyond $\sim 300$ Mpc. Overall, the time window for follow-up searches is shortest for distant BH-NS mergers and longest for nearby BNS mergers.
5 Conclusion
@@
Our analysis specifically explored four distinct BNS merger scenarios categorized by remnant mass and post-merger evolutionary pathways: stable neutron star formation (BNS-I), delayed black hole collapse following a supramassive neutron star phase (BNS-II), rapid black hole formation post-hypermassive neutron star phase (BNS-III), and immediate black hole formation (BNS-IV). Each scenario revealed characteristic X-ray emission patterns: notably, long-lived magnetar remnants (BNS-I and II) produce distinct, prolonged X-ray plateau phases powered by magnetar spin-down emissions, which significantly differentiate them from scenarios leading directly or rapidly to black hole formation.
Furthermore, we highlighted how viewing geometry dramatically influences observed X-ray signatures, classifying observation zones into jet-aligned (brightest emission), free (moderately dimmed emission), and trapped (strongly dimmed or delayed emission due to ejecta opacity) zones. Such geometrical considerations are useful for interpreting X-ray data from multi-messenger observations, especially in conjunction with gravitational wave detections.
@@
The code used to compute prompt X-ray emission from a relativistic jet and a magnetar-powered wind (see Sections 3.1 and 3.2) is publicly available under the MIT license on GitHub: github.com/cjules0609/PromptX. The code for modeling GRB afterglow emission (see Section 3.3 and Wang et al. 2025, in preparation) is publicly available under the MIT license on Github: github.com/YihanWangAstro/VegasAfterglow.

### v2

@@ IV.1.2 BNS-II (BNS $\rightarrow$ HMNS $\rightarrow$ SMNS $\rightarrow$ BH)
  • Jet zone: Bright prompt emission and decaying afterglow from the GRB jet. The extended wind emission becomes observable once the prompt emission sufficiently diminishes, and will decay quickly (prescribed by high-latitude emission effects) when the remnant collapses. See left panel of Fig. 11.
  • Free zone: Faint prompt emission and rising afterglow from the GRB jet. The extended X-ray wind emission becomes observable once the prompt emission sufficiently diminishes—in some cases the prompt emission will never exceed the wind emission and only a plateau will be observed. The plateau will decay quickly (prescribed by high-latitude emission effects) when the remnant collapses. See middle panel of Fig. 11.
  • Trapped zone: Extremely faint prompt emission and dim late-time rising afterglow from the GRB jet. If $t_{\rm thin}>t_{c}$, the extended wind emission will never be directly observable. See right panel of Fig. 11.
@@
The BNS-II scenario may not be conclusively ruled out for GRB 170817A due to a lack of X-ray coverage at very early times caused by localization delays, large positional uncertainties, and Earth occultation. This underscores the importance of continued advances in gravitational-wave localization pipelines and the development of wide-field and rapid-response X-ray detectors to enable prompt follow-up of future BNS mergers.
IV.1.3 BNS-III (BNS $\rightarrow$ HMNS $\rightarrow$ BH)
Physically, one would expect typical prompt and afterglow X-rays from the GRB jet without significant energy injection. The kilonova afterglow may be detectable in X-rays at much later times.
@@ IV.1.3 BNS-III (BNS $\rightarrow$ HMNS $\rightarrow$ BH)
The BNS-III scenario can naturally reproduce the light curves of GRB 170817A and does not necessitate a long-lived central engine.
IV.1.4 BNS-IV (BNS $\rightarrow$ BH)
Remnants with masses much larger than the TOV mass limit will be unable to withstand collapse into a BH for an appreciable amount of time. A jet may still be launched if there is matter outside the event horizon forming an accretion disk, and the associated light curve would likely resemble that of a BNS-III (Fig. 12)
As with the BNS-III scenario, the BNS-IV scenario also naturally reproduces the light curves of GRB 170817A.
IV.2 BH-NS Mergers
X-ray counterparts are only expected from BH-NS mergers that tidally disrupt the NS (BH-NS w/ TD). Physically, one might expect prompt and afterglow X-rays from a relativistic GRB jet. Compared to jets launched from BNS mergers, jets launched from BH-NS mergers w/ TD are likely less collimated (Bromberg et al., 2011; Duffell et al., 2015; Ruiz et al., 2018; Lazzati and Perna, 2019). Additionally, the dynamical ejecta should be directed primarily in the equatorial direction (Kyutoku et al., 2015; Kawaguchi et al., 2015, 2016), which affects the peak flux and time of X-ray signatures from BH-NS mergers (Sadeh, 2024). Therefore, we adopt $\theta_{c}=10^{\circ}$ and $\theta_{\rm cut}=40^{\circ}$ for a BH-NS w/ TD jet and adjust $E_{\rm iso}$ such that the beaming-corrected energy $E_{\gamma}=(1-\cos\theta_{c})\,E_{\gamma,\rm iso}$ remains constant. The BH-NS w/ TD jet, while less collimated than its BNS counterpart, has a shallower energy and Lorentz factor profile and larger cutoff angle, thus appearing brighter at large viewing angles.
@@ IV.2 BH-NS Mergers
Due to very strong evidence that GRB 170817A was associated with a BNS merger event—including the GW170817 signal (Abbott et al., 2017a), associated kilonova Cowperthwaite et al. (2017), redshift and localization, and our model parameters for BH-NS mergers—we definitively rule out the BH-NS model for GRB 170817A.
IV.3 The case of GW170817
The EM counterparts of GW170817, including both prompt and afterglow emission of GRB 170817A, can be consistently interpreted within the framework presented here. However, since the earliest X-ray coverage of the GW170817 final localization region were not until $10^{4}\,\rm s$ trigger, it remains unclear whether this BNS merger was a BNS-II or BNS-III/BNS-IV. Prompt X-ray detections of NS mergers by a rapid-slewing or wide-field X-ray detector can identify the presence of a long-lived central engine and distinguish between these scenarios.
IV.4 X-ray Counterpart Searches to Neutron Star Mergers
The search for EM counterparts of NS mergers detected by the LIGO-Virgo-KAGRA collaboration (LVK) is critical for advancing understanding of both the merger process and the Type I GRBs. The joint detection of GW170817 and GRB 170817A has opened a new era of multimessenger astronomy, demonstrating the power of combining GW and EM observations.
@@ IV.4 X-ray Counterpart Searches to Neutron Star Mergers
The newest generation of wide-field X-ray telescopes, including the Einstein Probe (EP; Yuan et al. (2025)) and the Space-based multi-band astronomical Variable Objects Monitor (SVOM; Wei et al. (2016)), boasts rapid response and dedicated sensitivity in the soft X-rays. EP’s Wide-field X-ray Telescope (WXT), with its wide field-of-view and high sensitivity in the soft X-ray band (0.50.544 keV), is designed to promptly detect and localize soft X-ray transients across broad sky regions. Its on-board Follow-up X-ray Telescope (FXT), allows deep follow-up observations in X-rays. SVOM offers complementary multi-wavelength capabilities through its diverse payload: the wide-field ECLAIR ($4-150\rm\,keV$) enables prompt detections, while the Microchannel X-ray Telescope (MXT; $0.2-10\rm\,keV$) facilitates rapid X-ray follow-up. Additionally, SVOM is equipped with rapid slewing capability and is integrated with ground-based follow-up networks, enhancing its effectiveness in capturing transient events.
@@
Fig. 6 shows the peak X-ray luminosity of prompt emission from NS mergers with EM counterparts observed at different viewing angles, for several jet opening angles and assuming a constant beaming-corrected energy $E_{\gamma,\rm tot}$. For BNS mergers, the jet opening angle is typically narrow, with $\theta_{c}\sim 3^{\circ}$ for GRB 170817A, whereas for BH-NS mergers, the opening angle is likely larger, $\theta_{c}\sim 10^{\circ}$ (Bromberg et al., 2011; Duffell et al., 2015; Ruiz et al., 2018; Lazzati and Perna, 2019). The true jet opening angle derived for GRB170817A remains unclear, with differing results from various methods (Gottlieb et al., 2018; Mooley et al., 2018a; Troja et al., 2019a; Ghirlanda et al., 2019; Ryan et al., 2024). This work is generally consistent with models invoking a single-component narrow GRB jet.
Figure 14: X-ray afterglow peak luminosity (left panel) and peak time (middle panel) as functions of viewing angle for several jet opening angles and assuming a constant beaming-corrected energy $E_{\gamma,\rm tot}$. The right panel shows the correlation between peak luminosity and peak time. The beaming-corrected energy is fixed at $10^{51}$ erg. The red star represents the best-fit values from our MCMC simulations for the X-ray afterglow peak luminosity and time for GRB 170817A. X-ray data was retrieved from Troja et al. (2020).
@@
With constraints from prompt X-ray emission, one may estimate the X-ray afterglow peak time and luminosity. Fig. 14 shows the X-ray afterglow peak luminosity (left panel) and peak time (middle panel) as functions of the viewing angle $\theta_{v}$ for several jet opening angles and assuming a constant beaming-corrected energy $E_{\gamma,\rm tot}$. With the inferred luminosity distance from LVK, observers may estimate the detection horizon for X-ray counterparts of NS mergers based on their detector’s sensitivity and time since the GW trigger (Fig. 16).
Figure 16: Detection horizon for X-ray (0.3-10 keV) afterglows from BNS mergers as a function of time, for given detector sensitivities. The model parameters are listed in Tab. 2.
@@
  • Prompt X-ray Analysis: Data from X-ray detectors that were covering the localization region during the trigger time should be examined for coincident events. If no events are found, upper limits can help constrain the prompt X-ray emission.
  • Geometric Constraints: Constraints on $\theta_{c}$, $\theta_{v}$, and $\theta_{\rm cut}$ for a GRB jet with beaming-corrected energy $E_{\gamma,\rm tot}$ corresponding to GRB 170817A can be derived from X-ray detections and upper limits using the model described in Section III.1 and illustrated in Fig. 6.
  • X-ray Afterglow Predictions: Given a fixed set of model parameters and geometric constraints from prompt emission, the X-ray afterglow characteristics (e.g., peak time and flux) can be estimated (Fig. 14). [^4] We emphasize that these predictions are highly sensitive to the chosen physical parameters and are only representative of bursts with properties similar to those assumed in the model. In this work, we adopt parameters representative of GRB 170817A, so the resulting predictions are specific to GW170817/GRB 170817A-like events.
@@
  • Coordinated Follow-up: X-ray follow-up observations can be planned based on the predicted afterglow properties. Using Fig. 16, one can estimate the optimal time window for detecting an X-ray counterpart of an NS merger with known luminosity distance, given one’s detector sensitivity. It suggests that prompt follow-up observations are very important. In case of non-detection at a certain epoch, one can estimate the luminosity distance horizon beyond which continued follow-up may no longer yield detections.
Additional detections of NS mergers with X-ray counterparts will be a major success of multimessenger astronomy.
Fig. 16 shows the detection horizon of a BNS-merger GRB afterglow with luminosity comparable to GRB 170817A. Here, we focus on the afterglow as it sets the long-term window for coordinated follow-up observations and, through its properties, provides constraints on the prompt emission and central engine activity that cannot be obtained from short-lived signals alone.
At any given epoch, there exists a range of viewing angles for which the afterglow peak emission remains above a given detector’s flux threshold. The detection horizon curves therefore represent the largest distance at which an off-axis afterglow exactly peaks at the threshold at a given time. Equivalently, for a fixed source distance, the afterglow would remain potentially detectable until the horizon curve drops below that distance; this includes cases where the off-axis afterglow is still rising and its peak has not yet occurred. For example, if EP-FXT (with sensitivity of about $10^{-14}\ {\rm erg/s/cm^{2}}$) does not detect an X-ray counterpart of a BNS merger like GW170817/GRB 170817A within one day, then follow-up observations may no longer yield detections if the source is further than $\sim 300$ Mpc. Equivalently, for a source with luminosity distance $D_{L}\sim 140\rm\,Mpc$ (e.g., obtained from a GW trigger as in GW170817) EP-FXT could successfully detect an X-ray counterpart within $\sim 100$ days.
The expected distribution of intrinsic events across viewing zones (jet, free, and trapped) is proportional to the solid angle subtended by each zone, $\Omega=4\pi\int_{\theta_{1}}^{\theta_{2}}\sin\theta\,d\theta=4\pi(\cos\theta_{2}-\cos\theta_{1})$, because observers are expected to be distributed isotropically. Therefore, the fractions of events observed in each zone can be expressed as
$f_{\rm jet}$ $=\frac{\Omega_{\rm jet}}{4\pi}=1-\cos\theta_{\rm c}$ (38)
$f_{\rm free}$ $=\frac{\Omega_{\rm free}}{4\pi}=\cos\theta_{\rm c}-\cos\theta_{\rm cut}$ (39)
$f_{\rm trapped}$ $=\frac{\Omega_{\rm trapped}}{4\pi}=\cos\theta_{\rm cut}.$ (40)
For BNS mergers ($\theta_{c}=3^{\circ}$ and $\theta_{\rm cut}=20^{\circ}$), $_{\rm jet}=0.0014, ff$$_{\rm free}=0.0589$, and $f_{\rm trapped}=0.9397$. This yields an intrinsic jet : free : trapped zone ratio of $\approx 1:42:671$. For BHNS mergers ($\theta_{c}=10^{\circ}$ and $\theta_{\rm cut}=40^{\circ}$), $_{\rm jet}=0.0152, ff$$_{\rm free}=0.2188$, and $f_{\rm trapped}=0.7660$. This yields an intrinsic jet:free:trapped zone ratio of $\approx 1:14:50$. We note, however, the observed jet:free:trapped zone ratios could be much different from these intrinsic values. This is because given a fixed detector sensitivity, the chance of detecting faint emission in the trapped or even free zones is much lower than that of detecting emission from the jet zone. A full analysis of detectability effects is beyond the scope of this paper.
V Conclusion
@@
Our analysis specifically explored four distinct BNS merger scenarios categorized by remnant mass and post-merger evolutionary pathways: stable neutron star formation (BNS-I), delayed black hole collapse following a supramassive neutron star phase (BNS-II), rapid black hole formation post-hypermassive neutron star phase (BNS-III), and prompt black hole formation (BNS-IV). Each scenario revealed characteristic X-ray emission patterns: notably, long-lived magnetar remnants (BNS-I and II) produce distinct, prolonged X-ray plateau phases powered by magnetar spin-down emissions, which significantly differentiate them from scenarios leading directly or rapidly to black hole formation.
Furthermore, we highlighted how viewing geometry dramatically influences observed X-ray signatures, classifying observation zones into jet-aligned (brightest emission), free (moderately dimmed emission), and trapped (strongly dimmed or delayed emission due to ejecta opacity) zones. Such geometrical considerations are useful for interpreting X-ray data from multimessenger observations, especially in conjunction with gravitational wave detections.
We performed a case study of GRB 170817A using MCMC simulations on multi-wavelength data, and showed that the prompt and afterglow frameworks developed in this paper are self-consistent and in agreement with similar models in the literature.
@@
The code used to compute prompt X-ray emission from a relativistic jet and a magnetar-powered wind (Sections III.1 and III.2) is publicly available under the MIT license on GitHub (github.com/cjules0609/PromptX) and archived in Zenodo (Chen et al., 2025). The code for modeling GRB afterglow emission (Section III.3) is publicly available under the BSD 3-clause license on Github: github.com/YihanWangAstro/VegasAfterglow (Wang et al., 2025).